OpenStudy (anonymous):

There is a circle with the center of (4,-1) and a radius of 7 the formula is (x-h)^2+ (y-k)^2=r^2 i think i got it but im not sure so i need a second opinion

OpenStudy (anonymous):

(x-4)^2+(y+1)^2=49

OpenStudy (anonymous):

hey thanks

OpenStudy (e.mccormick):

In \((x-h)^2+ (y-k)^2=r^2\) the center is the point \((h,k)\). So what alfonsopadilla did was look at what you were given: "center of (4,-1)" means \((h,k) = (4,-1)\) which in turn means: \(h=4\) \(k=-1\) "radius of 7" means \(r=7\) Then all he did was substitute in.